4. Determine the activation energy for the KI-catalysed decomposition of hydrogen peroxide. Things to consider: a) What variables do you need to determine the activation energy of a reaction? b) Which variable(s) would be dependent? Which would be c) How will you measure them? d) How many measurements do you need? e) How can you distribute the measurements among the group? independent?
Answer #1

As decomposition reaction is exothermic reaction where energy is released in the form of heat and this is used to determine the rate of a reaction.   

Energy released can easily be quantified than molecular concentration changes because change in temperature is easy to calculate using thermometers.

Or it can also calculated by change in concentration of hydrogen peroxide or volume or pressure of oxygen produced during decomposition.

But determination of rate of reaction at different temperature can easily calculated by a single run using simple calorimeter, determining temperature and time where activation energy (Ea) can be calculated by Arrhenius plot. (Natural logarithm of rate constant verses inverse of temperature will have a slope, m = -Ea/R.

Activation energy is independent of temperature and concentration of reactants and catalysts.

Arrehenius equation : This expresses the temperature dependence of the rate constant, k = Ae-(Ea/RT) where A is proportionate constant, R is ideal gas constant and T is temperature in kelvin.

Rate of reaction for KI catalysed decomposition of hydrogen peroxide :

   R = k [I- ]a [H2O2]b and k = Ae-(Ea/RT)  

   Rate = A exp (-Ea/RT) [I- ]a [H2O2]b

In (Rate) = (-Ea/RT) + In [I- ]a [H2O2]b

In (Rate) is dependent variable on Y -axis and Inverse of temperature (1/T) is independent variable on X-axis with slope m = -Ea/R.

Several measurements at differing concentration of iodide and plot of In (Rate/Rateo) as a function of In (I-/I-O) with slope a

In (Rate/Rateo) = a In (I-/I-O)   

Three Test tubes : Test tube 1 : Keep as reference

Test tube 2 : Add a drops of starch solution

Test tube 3 : Add a drops of sodium sulphate solution and add drop of starch solution

Experiment is repeated at room temperature, near 30, 35, 40, 45, and 50 C (Five total measurements per group)

Reaction rates can be evaluated by a) In (I-/I-O) - independent variable Vs   In (Rate/Rateo) -dependent variable

b) In (H2O2/H2O2O) - independent variable Vs   In (Rate/Rateo) -dependent variable

c) [I-]a - independent variable Vs Rate - dependent variable

d) Moles of H2O2 /sec independent variable Vs Rate - dependent variable

After measuring either of the above measurements, Other graph is plotted between natural logarithm of reaction time as a function of 1/T to determine the activation energy (Ea) of rate coefficient.